OpenStudy (anonymous):

Question:Rationalize the denominator of 3over 6-√5 Answers: 6+3√5 over 41 18+√5 over 31 18+3√5 over 31 18+3√5 over 41

OpenStudy (angela210793):

3(6+sqr5)/36-5=18+3sqr5/31=18+√5 over 31